Determine for which positive integer values of n, 3n^3+2≤n^4 and prove your claim by mathematical induction.


I was able to determine n must be greater or equal to 4. But I can't seem to get through the inductive step.


1
Expert's answer
2021-11-14T16:26:06-0500

Let "f(x)=3x^3+2-x^4, x>0"


"f'(x)=9x^2-4x^3"

Find the critical number(s)


"f'(x)=0=>9x^2-4x^3=0"

"x^2(9-4x)=0"

"x_1=0, x_2=2.25"

If "x<2.25, f'(x)>0, f(x)" increases.

If "x>2.25, f'(x)<0, f(x)" decreases.



"f(1)=3+2-1=4>0"

"f(2)=24+2-16=10>0"

"f(3)=81+2-81=2>0"

Therefore "3n^3+2\leq n^4" for "n\geq4".


Let "P(n)" be the proposition that "3n^3+2\leq n^4, n\geq4."

Basis Step

"P(4)" is true, because "3(4)^3+2=194<256=(4)^4."

Inductive Step

We assume that "3k^3+2\leq k^4." Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"3(k+1)^3+2\leq (k+1)^4"

We have


"3(k+1)^3+2=3k^3+9k^2+9k+3+2"

"\leq k^4+(4k)k^2+(6k)k+(4k+1)=(k+1)^4, k\geq4"

Hence "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for integers "n\geq4". That is, we have proven that


"3n^3+2\leq n^4, n\geq4."